A retailer estimated that the probability that the practice of basic commodity will increase, decrease, or
remain unchanged during a given month is 0.35, 0.10, and 0.55 respectively. What is the probability
that the price of the commodity will change during the month?

Found 2 solutions by greenestamps, ikleyn:
Answer by greenestamps(13200)



P(change) = P(increase OR decrease) = P(increase) + P(decrease) = 0.35+0.10 = 0.45.

OR...

P(change) = 1 - P(unchanged) = 1-0.55 = 0.45.

Answer by ikleyn(52781)

.

    P(change) = 1 - 0.55 = 0.45.


It is the same as to say


    P(change) = 0.35 + 0.10 = 0.45.



The answer is the same independently on the way how you calculate it.